OpenStudy (anonymous):

Predict the next number in this sequence: 1, 4, 9, 16, ... A. 25 B. 32 C. 23

OpenStudy (anonymous):

a

OpenStudy (anonymous):

it's a geometric sequence

OpenStudy (anonymous):

if you look at each number, you could square root each of them

OpenStudy (anonymous):

so the next would be 25, 36, 49, 64, 81, 100, 121, 144, 169, and so on

OpenStudy (anonymous):

can you show me an example on how you got 25

OpenStudy (anonymous):

4-1 is 3, so add 2 to three then add that to 4

OpenStudy (anonymous):

u get 9

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

9-4 is 5. add 2 then get 7. add to 9 and get 16 and notice that you could square root all of the numbers

OpenStudy (anonymous):

16-9 is 7, +2 is 9. 16+9 is 25

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

the easiest way to do it is to square the numbers

OpenStudy (zarkon):

\[1^2,2^2,3^2,4^2,5^2\]

OpenStudy (anonymous):

next question

OpenStudy (anonymous):

1^2=1 2^2=4 3^2=9 and so on

OpenStudy (anonymous):

my medal?

OpenStudy (anonymous):

The answer is 42.